During the latest year, XYZ Corporation has total sales of $500,000, net income of 30,000, and its year-end total assets were $250,000. The firm's total debt to total assets ratio was 50%. What is firm's return on equity (ROE)? (Please show all steps used to get the answer)
Answer #1

Debt to total assest ratio = 50%

Total debt = 0.50*250000 = 125000

Shareholders equity = total liabilities(total assests) -total debt

= 250000-125000 = 125000

Return on equity = Net income/Shareholders Equity

= 30000/125000 * 100

= 24%